# The Value of Context: Human versus Black Box Evaluators

### v1

@@ 2.4 Model Uncertainty
The type is equal to the value of nonstandard covariate $x_{I}$, where the index $I$ is drawn uniformly at random from $\mathcal{N}$.
@@
The value of $f(\mathbb{x}_{n})$ is (independently) drawn from a uniform distribution on $[1,2]$ if $x_{s+1}=1$, and (independently) drawn from a uniform distribution on $[0,1]$ if $x_{s+1}=0$.
We view Assumption 1 as a useful conceptual benchmark, and later explore how far our main results generalize under various relaxations of this assumption (Section 5). While our assumption of no prior knowledge about the role of nonstandard covariates is strong, by definition the nonstandard covariates are precisely the covariates for which there is little historical data. For example, while it may be well understood that a higher GPA correlates with higher on-the-job ability, a large number of social media followers could potentially be a positive or negative signal. Indeed, if it were understood that some nonstandard covariate was very predictive—for example, that having many social media followers was a strong indication of on-the-job success—we would expect this nonstandard covariate to become a standard covariate, and thus queried both by the human and black box evaluator.
@@
These are clearly very conservative criteria for what it means to prefer the human or the black box. In practice, we would expect that the set of revealed covariates $H$ to be intermediate to the two cases considered in Definition 2.5, i.e., that $H$ neither maximizes nor minimizes the agent’s payoffs. [^21] Angelova et al. (2022) provide evidence that some judges condition on irrelevant defendant covariates when predicting misconduct rates. But if we can conclude either that the agent prefers the black box evaluator or the human evaluator according to Definition 2.5, then the same conclusion would hold for these more realistic models of $H$.
@@ 3.1 Example
Suppose there are two covariates $x_{1}$ and $x_{2}$, both nonstandard. For each covariate vector $\mathbb{x}\in\{0,1\}^{2}$, define the random variable $Y_{\mathbb{x}}=f(\mathbb{x})$, where the randomness is in the realization of $f$.
@@ 3.2 The Expected Value of Context
Our main result says that for every agent, the expected value of context vanishes as $n$ grows large.
@@
Suppose Assumption 1 holds. Then for every covariate vector $\mathbb{x}\in\{0,1\}^{\infty}$,
$V(n,\mathbb{x}_{n})=O\left(\frac{\sqrt{n}}{2^{(1-\alpha_{h})n}}\right)$ (3.2)
Hence for any $\alpha_{h}<1$, the expected value of context vanishes to zero as $n$ grows large, i.e.,
@@
Thus although the value of context may be substantial for certain type functions (such as in Example 5), it does not matter on average across these functions when the agent’s prior satisfies Assumption 1. This also implies that provision of context does not “typically” matter; that is, the probability that the agent gains substantially from targeted information acquisition is small.
@@
This proof sketch also clarifies the role of Assumption 1. As we show in Section 5.3, the statement of the theorem extends so long as the evaluator’s posterior expectation $Z_{k}$ concentrates on its expectation sufficiently quickly as $n$ grows large. Roughly speaking, this means that the informativeness of any specific set of covariates is decreasing in the total number of covariates. Thus neither the precise symmetry imposed by Assumption 1, nor even the use of Bayesian updating, is critical for a result like Theorem 3.1 to hold.
@@
In Section 5 we explore several relaxations of Assumption 1: The first of these relaxations supposes that there is a “low-dimensional” set of covariates that predictive of the agent’s type, while the remaining covariates are irrelevant. The second relaxation supposes that there is a subset of nonstandard covariates whose effects are known. We formalize these extensions of our main model and examine the extent to which Theorem 3.1 extends.
@@ 3.3 Human versus Black Box
The agent’s expected utility can be written as $\mathbb{E}[\phi(\hat{y})]$ for some twice continuously differentiable function $\phi$ [^23] Restricting to utility functions that depend on a posterior mean is a common assumption in the literature on information design, see e.g., Kamenica and Gentzkow (2011), Frankel (2014) and Dworczak and Martini (2019).
@@
Suppose Assumptions 1 and 2 hold.
@@
    If $\phi$ is strictly convex, then there exists $N$ sufficiently large that the agent prefers the black box evaluator for all $n\geq N$.
@@
    If $\phi$ is strictly concave, then there exists $N$ sufficiently large that the agent prefers the human evaluator for all $n\geq N$.
Consider first the case of convex $\phi$ (Part (a)), corresponding to a preference for more accurate evaluations. [^24] Consider any two sets of covariates $A\subset A^{\prime}$ and let $\hat{y}_{A}$, $\hat{y}_{A^{\prime}}$ be the corresponding posterior expectations. The distribution of $\hat{y}_{A^{\prime}}$ (i.e., the posterior expectation that conditions on more information) is a mean-preserving spread of the distribution of $\hat{y}_{A}$. When $\phi$ is convex, the former leads to a higher expected utility. Such an agent “prefers more accurate evaluations” in the sense that giving the evaluator better information (in the standard Blackwell sense) leads to an improvement in the agent’s expected utility. Such an agent prefers for the evaluation to be based on more information (advantaging Black Box), but also prefers for the evaluation to be based on more relevant covariates (advantaging Human). We use the rate of convergence demonstrated in Theorem 3.1 (see (3.2)) to show that what eventually dominates is how many covariates the evaluators observe, not how they are selected; for an agent who prefers accuracy, this favors the Black Box.
Part (b) of Theorem 3.2 says that if instead $\phi$ is concave (corresponding to an agent who prefers inaccurate evaluations) then the agent should eventually prefer the human evaluator. We conclude this section with example decision problems that induce utility functions satisfying the conditions of either part of the theorem.
@@
Suppose the agent’s type is $y\in\{0,1\}$, and the evaluator chooses an action $a$ based on the observed covariates. The evaluator and agent share the utility function $-\mathbb{E}[(a-y)^{2}]$. The evaluator’s optimal action is $a=\hat{y}$, and the agent’s expected payoff given this action is $\mathbb{E}\left[\hat{y}^{2}-\hat{y}\right]$. So Part (b) of Theorem 3.2 says that the agent eventually prefers evaluation by the black box evaluator. [^25] Although the conditions of Theorem 3.2 are no longer met when $y$ is not binary, we show in Appendix A.6 that the conclusion of Part (b) of Theorem 3.2 generalizes for arbitrary $y$ given the mean-squared error payoff function described in this example.
@@ 4.1 Max value of context across agents
So far we have studied the the expected value of context for a single agent. If we instead ask whether the firm should use human or algorithmic evaluation—for example, whether a hospital should automate diagnoses or rely on doctor evaluations—various other statistics may also be relevant. For example, it may matter whether the value of context is large for any agent in the population (e.g., because a lawsuit regarding algorithmic error may be brought on the basis of harm to any specific individual (Jha, 2020)). We thus study the expected maximum value of context, as defined below.
@@
[…]
@@
[…]
@@ 4.2 Strategic Disclosure
Suppose Assumption 1 holds. Then for all $n$,
[…]
@@
5 Relaxing the Symmetry Assumption
As shown in Example 11, our main results can fail if the assumption of symmetric uncertainty over the role of the nonstandard covariate values (Assumption 1) is broken. We now propose three particular variations on Assumption 1 and explore the extent to which our main result extends. In Section 5.1, we suppose that it is known ex-ante that some $r_{n}$ covariates are relevant, while the remaining $n-r_{n}$ are not, so that even as $n$ grows to infinity, the effective number of covariates potentially grows more slowly. In Section 5.2 we allow the agent to have prior knowledge about the role of certain nonstandard covariates. Finally, Section 5.3 provides an abstract condition on the learning environment under which our main results hold, which requires the evaluator’s uncertainty about the agent’s type to grow sufficiently fast in $n$.
5.1 Irrelevant covariates
@@
To model this, we suppose there is a sequence of sets of relevant covariates $(R_{1},R_{2},\dots)$ such that each $R_{n}$ includes the standard covariates in $\mathcal{S}$ and is of size $s+r_{n}$, where $r_{n}=\lfloor\alpha_{r}\cdot n\rfloor$ is the (known) number of relevant nonstandard covariates. The irrelevance of the remaining covariates is reflected in the following assumption, which says that, holding fixed the values of the relevant covariates, the values of the irrelevant covariates do not change the type.
@@
For every $x_{R_{n}}\in\{0,1\}^{s+r_{n}}$ and $x_{-R_{n}},x^{\prime}_{-R_{n}}\in\{0,1\}^{n-s-r_{n}}$,
$f(x_{R_{n}},x_{-R_{n}})=f(x_{R_{n}},x^{\prime}_{-R_{n}})$
We then modify Assumption 1 to impose symmetry only over realizations of the relevant covariates.
Assumption 4.
@@
Suppose Assumption 1 holds and $\alpha_{h}<\alpha_{r}$. Then for every covariate vector $\mathbb{x}\in\{0,1\}^{\infty}$ the expected value of context vanishes to zero as $n$ grows large.
@@
This result suggests that whether human or black box evaluation is more appropriate should be determined in part based on whether the available signal is concentrated in a small number of covariates (favoring the human evaluator) or spread out across a large number of covariates (favoring the black box evaluator). The same application may transition between these regimes over time. For example, in a medical setting where black box diagnosis is highly accurate based on non-interpretable features of an image scan, it may not be possible to communicate sufficient information via any small number of covariates. But if the predictive features of the image are subsequently better understood and defined, then it may be that a small set of (newly defined) features does eventually capture all of the signal content, and can be fully disclosed in a coversation.
5.2 Known effect of certain covariates
@@
Specifically suppose there is a set $K\subseteq\{1,\dots,n\}$ of covariate indices whose effects are known. The set $K$ includes the standard covariates, but possibly also includes some nonstandard covariates. We weaken Assumption 1 to the following:
Assumption 5.
Fix any realization of the covariates $\mathbb{x}_{K}=(x_{i})_{i\in K}$. Then there is an infinitely exchangeable sequence $(\widetilde{Y}_{1},\widetilde{Y}_{2},\dots)$ such that for every $n\in\mathbb{N}$, the sequence
$\left(Y_{\mathbb{x}_{K},\mathbb{x}_{-K}}:\mathbb{x}_{-K}\in\{0,1\}^{n-|K|}\right)$
has the same distribution as $(\widetilde{Y}_{1},\widetilde{Y}_{2},\dots)$.
This assumption imposes exchangeability only over the nonstandard covariates whose effects are not ex-ante known. Clearly if $K$ is a strict superset of $\mathcal{S}$, then the expected value of context need not vanish. A simple example is the following.

### v2

@@ 2.4 Model Uncertainty
The type is equal to the value of the nonstandard covariate $x_{I}$, where the index $I$ is drawn uniformly at random from $\mathcal{N}$. Then Assumption 1 fails. [^19] Suppose $n=2$, and both covariates are nonstandard. Then under the agent’s prior, $f\in\left\{\hat{f},\tilde{f}\right\}$ where $\hat{f}(1,1)=\hat{f}(1,0)=1$ while $\hat{f}(0,1)=\hat{f}(0,0)=0$, and $\tilde{f}(1,1)=\tilde{f}(0,1)=1$ while $\tilde{f}(1,0)=\tilde{f}(0,0)=0$. So the agent knows with certainty that $f(1,1)=1$ but $f(0,0)=0$, in violation of Assumption 1. This example is consistent with exchangeability in the labels of the nonstandard covariates, but not with exchangeability in their realizations.
@@
The value of $f(\mathbb{x}_{n})$ is (independently) drawn from a uniform distribution on $[1,2]$ if $x_{s+1}=1$, and (independently) drawn from a uniform distribution on $[0,1]$ if $x_{s+1}=0$. Then Assumption 1 fails.
We view these two assumptions as useful conceptual benchmarks, but neither is necessary for our subsequent results. In Section 5, we explore how far our main results generalize under different relaxations of Assumptions 1 and 2.
2.5 Expected Value of Context
@@
These are clearly very conservative criteria for what it means to prefer the human or the black box. In practice, we would expect the set of revealed covariates $H$ to be intermediate to the two cases considered in Definition 2.3, i.e., that $H$ neither maximizes nor minimizes the agent’s payoffs. [^21] Angelova et al. (2022) provide evidence that some judges condition on irrelevant defendant covariates when predicting misconduct rates. But if we can conclude either that the agent prefers the black box evaluator or the human evaluator according to Definition 2.3, then the same conclusion would hold for these more realistic models of $H$.
@@ 3.1 Example
Suppose there are two covariates, $x_{1}$ and $x_{2}$, both nonstandard. For each covariate vector $\mathbb{x}\in\{0,1\}^{2}$, define the random variable $Y_{\mathbb{x}}=f(\mathbb{x})$, where the randomness is in the realization of $f$.
@@ 3.2 The Expected Value of Context
Our main result says that for every agent, the expected value of context (as defined in Definition 2.2) vanishes as $n$ grows large.
@@
Suppose Assumptions 1 and 2 hold. Then for every covariate vector $\mathbb{x}\in\{0,1\}^{\infty}$, the expected value of context vanishes to zero as $n$ grows large, i.e.,
@@
Thus although the value of context may be substantial for certain type functions (such as in Example 5), it does not matter on average across these functions when the agent’s prior satisfies Assumptions 1 and 2. This also implies that for sufficiently large $n$, the provision of context does not “typically” matter; that is, the probability that the agent gains substantially from targeted information acquisition is small.
@@
This proof sketch also clarifies the role of Assumption 1. As we show in Section 5.4, the statement of the theorem extends so long as the evaluator’s posterior expectation $Z_{k}$ concentrates on its expectation sufficiently quickly as $n$ grows large. Roughly speaking, this means that the informativeness of any specific set of covariates is decreasing in the total number of covariates. Thus the precise symmetry imposed by Assumption 1 is not critical for Theorem 3.1 to hold.
@@
In Section 5 we explore several relaxations of Assumptions 1 and 2. Our first relaxation of Assumption 1 supposes that there is a “low-dimensional” set of covariates that predictive of the agent’s type, while the remaining covariates are irrelevant. The second relaxation supposes that there is a subset of nonstandard covariates whose effects are known. We also consider a relaxation of Assumption 2 where the evaluator’s ability to predict $Y$ is increasing in the total number of covariates that define the type. We formalize these extensions of our main model and examine the extent to which Theorem 3.1 extends.
@@ 3.3 Human versus Black Box
The agent’s expected utility can be written as $\mathbb{E}[\phi(\hat{y})]$ for some twice continuously differentiable function $\phi$ [^23] Restricting to utility functions that depend on a posterior mean is a common assumption in the literature on information design, see e.g., Kamenica and Gentzkow (2011), Frankel (2014) and Dworczak and Martini (2019). Moreover, there exists $C<\infty$ such that
$\frac{\sup_{\hat{y}\in[-\overline{y},\overline{y}]}|\phi^{\prime}(\hat{y})|}{\inf_{\hat{y}\in[-\overline{y},\overline{y}]}|\phi^{\prime\prime}(\hat{y})|}<\frac{C}{2}.$
Roughly speaking, the numerator describes the sensitivity of the function $\phi$ to the evaluation, and the denominator describes the curvature of the function $\phi$. The assumption thus says that the curvature of the function must be sufficiently large relative to its slope. While there is no formal relationship, the LHS is evocative of the coefficient of absolute risk aversion of the function $\phi$ [^24] Recall that the coefficient of absolute risk aversion of the function $\phi$ is $-\frac{\phi^{\prime}(\hat{y})}{\phi^{\prime\prime}(\hat{y}))}$.
@@
Suppose Assumptions 1-3 hold, and let
$N=\min\left\{n\in\mathbb{R}_{+}:(\alpha_{b}-\alpha_{h})n-\frac{1}{2}\log_{2}(n)-1>\log_{2}(C)\right\}.$ (3.4)
Then:
@@
    If $\phi$ is strictly convex, the agent prefers the black box evaluator for all $n\geq N$.
@@
    If $\phi$ is strictly concave, the agent prefers the human evaluator for all $n\geq N$.
The comparisons in this theorem apply for reasonably small $N$. For example, let $C=100$, in which case the restriction in Assumption 3 is quite weak. Figure 1 fixes $\alpha_{h}=0.1$ and plots $N$ for different values of $\alpha_{b}$. If (say), the human evaluator observes 10% of covariates while Black Box observes 90%, then the comparisons in Theorem 3.2 hold for all $n\geq 14$.
Figure 1: Let $C=100$ and $\alpha_{h}=0.1$. Then the comparisons in Theorem 3.2 hold for all $n\geq N$ as depicted here.
The case of convex $\phi$ (Part (a)) corresponds to a preference for more accurate evaluations. [^25] Consider any two sets of covariates $A\subset A^{\prime}$ and let $\hat{y}_{A}$, $\hat{y}_{A^{\prime}}$ be the corresponding posterior expectations. The distribution of $\hat{y}_{A^{\prime}}$ (i.e., the posterior expectation that conditions on more information) is a mean-preserving spread of the distribution of $\hat{y}_{A}$. When $\phi$ is convex, the former leads to a higher expected utility. Such an agent “prefers more accurate evaluations” in the sense that giving the evaluator better information (in the standard Blackwell sense) leads to an improvement in the agent’s expected utility. Such an agent prefers for the evaluation to be based on more information (advantaging Black Box), but also prefers for the evaluation to be based on more relevant covariates (advantaging Human). We show that what eventually dominates is how many covariates the evaluators observe, not how they are selected; for an agent who prefers accuracy, this favors the Black Box.
Part (b) of Theorem 3.2 says that if instead $\phi$ is concave, then the agent should eventually prefer the human evaluator. We conclude this section with example decision problems that induce utility functions satisfying the conditions of either part of the theorem.
@@
Suppose the agent’s type is $y\in\{0,1\}$, and the evaluator chooses an action $a$ based on the observed covariates. The evaluator and agent share the utility function $-\mathbb{E}[(a-y)^{2}]$. The evaluator’s optimal action is $a=\hat{y}$, and the agent’s expected payoff given this action is
$\mathbb{E}\left[-(\hat{y}-y)^{2}\right]=\mathbb{E}\left[-\left(\hat{y}(1-\hat{y})^{2}+(1-\hat{y})(\hat{y})^{2}\right)\right]=\mathbb{E}\left[\phi(\hat{y})\right]$
where $\phi(\hat{y})=\hat{y}^{2}-\hat{y}$ is convex. So Part (a) of Theorem 3.2 says that the agent eventually prefers evaluation by the black box evaluator. Although the conditions of Theorem 3.2 are no longer met when $y$ is not binary, we show in Appendix A.7 that the conclusion of Part (a) of Theorem 3.2 generalizes for arbitrary $y$ given the mean-squared error payoff function described in this example.
@@ 4.1 Max value of context across agents
So far we have studied the the expected value of context for a single agent. If we instead ask whether the firm should use human or algorithmic evaluation—for example, whether a hospital should automate diagnoses or rely on doctor evaluations—other statistics may also be relevant. For example, it may matter whether the value of context is large for any agent in the population (e.g., because a lawsuit regarding algorithmic error may be brought on the basis of harm to any specific individual (Jha, 2020)). We thus study the expected maximum value of context, as defined below.
@@
$V^{\text{max}}(n)=\mathbb{E}\left[\max_{\mathbb{x}_{n}\in\{0,1\}^{n}}v(f,\mathbb{x}_{n})\right].$
@@
Suppose Suppose Assumptions 1 and 2 hold. Then the expected maximum value of context vanishes to zero as $n$ grows large, i.e., $\lim_{n\rightarrow\infty}\overline{V}^{\text{max}}(n)=0.$
@@ 4.2 Strategic Disclosure
Suppose Assumptions 1 and 2 hold. Then for all $n$,
$V^{D}(n)\leq V^{\text{max}}(n).$
@@
5 Relaxing our Assumptions on the Prior
As shown in Example 11, our main results can fail if the assumption of symmetric uncertainty over the role of the nonstandard covariate values (Assumption 1) is broken. We now propose two relaxations of Assumption 1 and one relaxation of Assumption 2, and explore the extent to which our main result extends. In Section 5.1, we suppose that it is known ex-ante that some $r_{n}$ covariates are relevant, while the remaining $n-r_{n}$ are not, so that even as $n$ grows to infinity, the effective number of covariates potentially grows more slowly. In Section 5.2 we allow the agent to have prior knowledge about the role of certain nonstandard covariates. In Section 5.3, we consider a model in which the predictability of the agent’s type is increasing in the total number of covariates. Finally, Section 5.4 provides an abstract condition on the learning environment under which our main results hold, which requires the evaluator’s uncertainty about the agent’s type to grow sufficiently fast in $n$.
5.1 Not all covariates are relevant
@@
To model this, we suppose there is a sequence of sets of relevant covariates $(R_{1},R_{2},\dots)$ such that each $R_{n}$ includes the standard covariates in $\mathcal{S}$ and is of size $s+r_{n}$, where $r_{n}=\lfloor\alpha_{r}\cdot n\rfloor$ is the (known) number of relevant nonstandard covariates. We will moreover without loss index these sets so that the relevant covariates are $x_{s+1},x_{s+2},\dots,x_{s+r_{n}}$. The irrelevance of the remaining covariates is reflected in the following assumption, which says that, holding fixed the values of the relevant covariates, the values of the irrelevant covariates do not change the type.
@@
There is a function $g(x_{1},\dots,x_{s+r_{n}})$ such that
$f(x_{1},\dots,x_{n})=g(x_{1},\dots,x_{s+r_{n}})$
for every $(x_{1},\dots,x_{n})\in\{0,1\}^{n}$.
We then modify Assumptions 1 and 2 to apply only to the relevant covariates.
Assumption 5 (Exchangeability).
For every realization of $(x_{1},\dots,x_{s})$, the sequence
$(\widetilde{Y}^{n}_{1},\dots,\widetilde{Y}^{n}_{2^{n}})\equiv(g(x_{1},\dots,x_{s},x_{s+1},\dots,x_{s+r_{n}}):(x_{s+1},\dots,x_{s+r_{n}})\in\{0,1\}^{r_{n}-s})$ (5.1)
is finitely exchangeable.
Assumption 6 (Constant Unpredictability of $Y$).
For every realization of the standard covariates $(x_{1},\dots,x_{s})$ and every pair $n^{\prime}>n$, the sequence $(\widetilde{Y}^{n}_{1},\dots,\widetilde{Y}_{2^{s+r_{n}}}^{n})$ and the truncated sequence $(\widetilde{Y}^{n^{\prime}}_{1},\dots,\widetilde{Y}^{n^{\prime}}_{2^{s+r_{n}}})$ are identical in distribution.
@@
Suppose Assumptions 5 and 6 hold, where $\alpha_{h}<\alpha_{r}$. Then for every covariate vector $\mathbb{x}\in\{0,1\}^{\infty}$ the expected value of context vanishes to zero as $n$ grows large.
@@
This result suggests that whether human or black box evaluation is more appropriate should be determined in part based on whether the available signal is concentrated in a small number of covariates (favoring the human evaluator) or spread out across a large number of covariates (favoring the black box evaluator). The same application may transition between these regimes over time. For example, in a medical setting where black box diagnosis is highly accurate based on non-interpretable features of an image scan, it may not be possible to communicate sufficient information via any small number of covariates. But if the predictive features of the image are subsequently better understood and defined, then it may be that a small set of (newly defined) features does eventually capture all of the signal content, and can be fully disclosed in a conversation.
5.2 Certain nonstandard covariates have known effects
@@
Specifically suppose there is a set $K\subseteq\{1,\dots,n\}$ of covariate indices whose effects are known. The set $K$ includes the standard covariates, but possibly also includes some nonstandard covariates. Without loss, we can index these $x_{1},\dots,x_{K}$. We weaken Assumption 1 to the following:
Assumption 7 (Exchangeability).
For every realization of the covariates $(x_{1},\dots,x_{K})$,
$(Y^{n}_{1},\dots,Y^{n}_{2^{n}})\equiv(f(x_{1},\dots,x_{K},x_{K+1},\dots,x_{n}):(x_{K+1},\dots,x_{n})\in\{0,1\}^{n-|K|})$ (5.2)
is finitely exchangeable.
This assumption imposes exchangeability only over the nonstandard covariates whose effects are not ex-ante known. Clearly if $K$ is a strict superset of $\mathcal{S}$, then the expected value of context need not vanish under Assumptions 2 and 7. A simple example is the following.
@@
5.3 Information accumulates in $n$
So far we have assumed that the predictability of $Y$ is constant in the number of covariates. This is not essential for our results. Suppose instead that
$y_{n}=f(x_{1},\dots,x_{n})+\varepsilon_{n}$
where $\varepsilon_{n}$ is a mean-zero random variable that is independent of the covariates $(x_{1},\dots,x_{n})$. This describes a setting in which the covariates are not sufficient to reveal the agent’s type, and there is a residual unknown.
Our previous results extend directly when the distribution of $\varepsilon_{n}$ is the same for all $n$. Another natural case is one in which $\text{Var}(\varepsilon_{n})$ decreases monotonically in $n$, with $\lim_{n\rightarrow\infty}\text{Var}(\varepsilon_{n})=0$; that is, in environments with a larger number of covariates, the agent’s type is more predictable. In Appendix B.5 we show that Theorem 3.1 directly extends. We also show that the comparisons in parts (a) and (b) of Theorem 3.2 hold for sufficiently large $n$, under the following assumption:
Assumption 8.
For each $n\in\mathbb{Z}_{+}$, let $\sigma_{\varepsilon,n}^{2}:=\text{Var}(\varepsilon_{n})$ and assume that $\frac{\varepsilon_{n}}{\sqrt{\sigma_{\varepsilon,n}^{2}}}$ admits a pdf, which we denote by $g_{n}$. The sequence $\{g_{n}(0)\}_{n}$ is bounded.
Loosely speaking, our comparisons in Theorem 3.2 continue to hold so long as the variance of $\varepsilon_{n}$ does not increase too fast in $n$.